The selling price of a couch at a furniture store is $35 more than twice the wholesale price. If the selling price is $475, what is the wholesale price?
Answer by richwmiller(17219)

sp=(2*wp)+35
sp=475
475=(2*wp)+35
440=2*wp
Finish the problem